# Insuring uninsurable income

### v15

@@ Abstract
We study dynamic mechanism design in a pure-exchange economy with privately observed idiosyncratic income. Classic hidden-income contracts attain constrained efficiency only at the cost of immiseration (Green 1987; Thomas–Worrall 1990). We propose a simple recursive mechanism—adapted from Marcet–Marimon (1992)—that shifts each income shock forward by one period, keeps promised utilities in a bounded set, and, under a transparent “moderate risk-aversion” condition, delivers sequential efficiency. In a stationary overlapping-generations setting, we further provide an explicit condition on the initial promise that ensures budget sustainability; early cohorts pre-fund intertemporal smoothing so that every cohort attains a higher expected lifetime utility than under autarky. Our analysis uses a single state (promised utility), closed-form transfers, and a Bellman verification.
@@ 1 Introduction
Private information can render idiosyncratic income effectively uninsurable. Since the seminal work of Green [1987] and Thomas and Worrall [1990], the constrained-efficient hidden-income contract trades insurance off against incentives and features a drift of average promised utilities toward a lower bound—immiseration—with cross-sectional dispersion rising over time. [^2] Earlier papers sometimes use immiserization[e.g., Phelan, 1998, Zhang, 2009]. Recent work with persistent private information, Bloedel et al. [2025], also sharpens the mechanism of immiseration and explains the balance of the cost of incentive provision and backloading of high-powered incentives. When agents are almost perfectly patient, the drift can vanish [Carrasco et al., 2019], but the knife-edge offers little comfort for moderate discounting.
@@
This paper adapts the Marcet–Marimon risk-shifting mechanism to an exchange economy in which agents cannot save. Promised utility is the sole state variable, so the model yields a closed-form characterization. If individual assets were allowed, one would require the set-valued approximation methods of Sleet and Yeltekin [2016]—a contrast confirmed by our closed-form example in Section 4.
The contribution of my paper is threefold. (i) A recursive, one-state dynamic mechanism that postpones shocks one period ahead and eliminates immiseration while preserving sequential incentive compatibility; (ii) a transparent sufficient condition for sequential efficiency under moderate risk aversion; (iii) in a stationary overlapping-generations society, an explicit inequality on the initial promise that suffices for budget sustainability, with a simple economic interpretation; early cohort pre-fund later transfers; the mechanism therefore delivers higher expected lifetime utility than autarky for all cohorts.
@@ 2.1 Environment
We consider a risk-neutral planner allocating transfers across a continuum of risk-averse individuals indexed by birth cohorts. Individuals survive to the next period with probability $\alpha\in(0,1)$ and discount at rate $r\in(0,1)$; their effective discount factor is $\beta=\alpha r$. In each period, a newborn cohort enters so that the cross-section is stationary. [^3] A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game. Each individual draws an i.i.d. income shock $e_{t}\in E:=\{e^{1},e^{2},\dots,e^{M}\}\subset\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$. The planner commits to a recursive transfer mechanism $\mathcal{M}$ mapping the current promised $\lambda_{t}>0$ into a transfer $\tau_{t}$ and next period’s promise $\lambda_{t+1}$. The single state variable is the promised utility $\overline{v}_{1}(\lambda)$, which is strictly increasing in $\lambda$.
@@ 2.2 Symmetric-information benchmark
where $\overline{C}>0$ is an exogenous upper bound on the present value of net subsidies, $\mathbb{E}_{0}$ is the conditional expectation given the information in period 0, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period. The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above.
@@
$J(\lambda_{t},t,\tau)=(1-\beta)\mathbb{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T}.$ (2)
@@ 2.3 Private information and mechanism design
Definition 1.
[Sequential incentive compatibility, Marcet and Marimon, 1992] Given the current promise $\lambda$, truthful reporting maximizes the continuation value. Equivalently,
@@
Definition 2.
[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be a sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequentially incentive compatible mechanism.
@@
Such a promise $\lambda^{\prime}$ that satisfies both (3) and (4) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$. By a feasible $\lambda^{\prime}$ we mean that the updated promise keeps the level of expected utility for the next period below the supremum of each individual’s utility function. That is, let $S$ be the supremum of each individual’s utility function, $S=\sup_{x\in\mathbb{R}_{+}}u(x)$. For given $\lambda>0$, if $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))<S$, define $\lambda^{\prime}$ such that
$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))\big{)}.$ (5)
Under the updating rule of $\lambda^{\prime}$ in (5), misreporting never raises the weighted sum $\lambda u(e+\tau^{*}(e|\lambda))+v_{2}(\lambda,e)$, hence truth-telling is optimal.
@@ 3 Risk-shifting mechanism
Definition 3.
(Risk-shifting mechanism) A mechanism $\mathcal{M}$ is a sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ and transfers $\tau_{t}(e_{t})=\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ such that $\lambda_{t+1}$ is given by (6). Within each period $t$, consumption is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda_{t}^{-1})$; schocks affect only $\lambda_{t+1}$.
@@
Interpretation. Arrow-Pratt risk aversion at the contract consumption level is smaller than a scaled income spread; the condition is mild when $\beta$ is close to one. Under Assumption 1, feasibility does not bind in the promise-update map, and the recursive rule (5) applies period by period.
@@
We could remove condition Assumption 1 by removing the assumption of bounded utility; dropping it would not violate sequential efficiency but would complicate the martingale proof. Boundedness is used to ensure that $\{\overline{v}_{1}(\lambda_{t})\}$ converges.
@@ 3.1 Individual level: bounded promised utilities
This subsection shows that the sequence of promised utilities forms a bounded martingale and therefore converges, so immiseration cannot occur.
@@
The promised utility process $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ is a bounded martingale and converges to an integrable random variable a.s. (no immiseration).
@@
Because promised utility is a martingale with bounded support, the mechanism is a fair intertemporal lottery: expected lifetime utility is preserved ex-ante, but its realisation drifts neither up nor down on average.
@@ 3.2 Planner’s budget and sustainability
Theorem 3.
@@
Proof: Second-order approximation shows decreasing differences. See Appendix C.
@@ 4 Numerical examples
An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Theorem 3 is given by the second-order polynomial approximation as follows.
@@
The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Theorem 3 holds with the second-order polynomial approximation if the initial promise $\lambda_{0}$ satisfies
@@
We illustrate the mechanism with a CRRA utility, a symmetric discrete income distribution, and the benchmark parameter $r=\alpha=0.93$.
@@
The values follow Corollary 1: the lower bound $\lambda_{CE}=1/u^{\prime}\bigl{(}E[u(e_{t})]\bigr{)}=49.608$ ensures higher lifetime utility than autarky across all cohorts, and the upper bound $133$ is set to satisfy (9).
@@
The series of sample means of transfer balances $\{g_{t}(\lambda_{0})\}_{t=0}^{1000}$ remains non-negative, and supports Theorem 3.
@@
Remark 3.
Classic hidden-income contracts [e.g. Green, 1987, Thomas and Worrall, 1990] spread shocks within each period: truth telling is rewarded by a higher spot transfer, so lifetime utilities diverge over time and immiseration follows. Under our mechanism, by contrast, the realized income $e_{t}$ only determines next period’s promise $\lambda_{t+1}$; consumption in period $t$ is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda^{-1}_{t})$ with probability $1$. The shock is therefore shifted one step forward—a transfer of risk to the future. This intertemporal insurance does make the average transfer profile $\{-\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ fall with age, threatening the planner’s long-run budget. Theorem 3 shows the problem can be neutralized by choosing an initial promise $\lambda_{0}$ that gives early cohorts negative net transfers, so later positive transfers are prefunded. Finally, note a key difference from Marcet and Marimon [1992]: in their growth model, investment returns are themselves the incentive, and the contract is bilateral and non-competitive, which can depress the manager’s utility below autarky. Our pure-exchange setting has no such external scale effect, so every cohort attains a higher lifetime utility than autarky.

### v16

@@ Abstract
We study dynamic mechanism design in a pure-exchange economy with privately observed idiosyncratic income. In the standard infinitely lived hidden-income benchmark of Green (1987) and Thomas–Worrall (1990), constrained-efficient allocations exhibit immiseration. We propose a simple recursive mechanism—adapted from Marcet–Marimon (1992)—that shifts each income shock forward by one period, keeps promised utilities in a bounded set, and, under a transparent “moderate risk-aversion” condition, delivers sequential efficiency. In a stationary overlapping-generations setting, we further show that under additional symmetry and curvature assumptions, a second-order approximation yields a sufficient condition for period-by-period budget balance; early cohorts pre-fund later transfers; for suitable initial promises, all cohorts are better off than under autarky. Our analysis uses a single state (promised utility), closed-form transfers, and a Bellman verification.
@@ 1 Introduction
Private information can render idiosyncratic income effectively uninsurable. Since the seminal work of Green [1987] and Thomas and Worrall [1990], the constrained-efficient hidden-income contract trades insurance off against incentives and features a drift of average promised utilities toward a lower bound—immiseration—with cross-sectional dispersion rising over time. [^2] Earlier papers sometimes use immiserization[e.g., Phelan, 1998, Zhang, 2009]. When agents are almost perfectly patient, the drift can vanish [Carrasco et al., 2019], but recent work with persistent private information, Bloedel et al. [2025], shows that immiseration remains robust beyond the i.i.d. benchmark.
@@
This paper adapts the Marcet–Marimon risk-shifting mechanism, that is, the current private shock changes tomorrow’s promised utility rather than today’s consumption, to an exchange economy in which agents cannot save. Promised utility is the sole state variable, so the model yields a closed-form characterization.
Relative to Green [1987] and Thomas and Worrall [1990], this paper changes three ingredients: agents have stochastic finite lifetimes in a stationary OLG environment, the mechanism is restricted to full within-period insurance, and incentives are provided through continuation value alone. The paper studies the consequences of that restricted class of mechanisms.
The contribution of this paper is threefold. (i) A recursive, one-state dynamic mechanism that postpones shocks one period ahead and eliminates immiseration while preserving sequential incentive compatibility; (ii) a transparent sufficient condition for sequential efficiency under moderate risk aversion; (iii) in a stationary overlapping-generations society, an explicit inequality on the initial promise that suffices for budget sustainability, with a simple economic interpretation; early cohorts pre-fund later transfers; the mechanism therefore delivers higher expected lifetime utility than autarky for all cohorts for suitable initial promises.
Unlike Fernandes and Phelan [2000], who enlarge the recursive state because serial correlation in private shocks destroys the one-dimensional promised-utility representation, this paper retains an i.i.d. hidden-income environment but identifies a mechanism in which current private shocks are absorbed by continuation promises. Under Assumption 1, this yields a simple one-state Bellman representation with full within-period insurance and bounded promised utility.
Farhi and Werning [2007] and Phelan [2006] reinterpret and extend the classic hidden-information benchmark by replacing an infinitely lived individual with a dynasty of altruistically linked one-period-lived agents. In that dynastic literature, the central issue is how welfare, or promised utility, is inherited across parent and child under the incentive logic of Green (1987) and Thomas–Worrall (1990). By contrast, this paper studies a stationary overlapping-generations environment with stochastic survival. Its mechanism provides incentives by shifting the current private shock into next period’s promised utility; under Assumption 1, this is consistent with full within-period insurance.
Demange [2002] studies Pareto properties of equilibrium risk sharing in stochastic OLG economies and identifies conditions under which adding a PAYG system at equilibrium is not Pareto-improving. By contrast, this paper studies a planner’s sequentially incentive-compatible transfer mechanism under hidden income. The mechanism is not a pure equilibrium PAYG reform: it uses continuation promises and, in the OLG extension, can involve transition pre-funding by early cohorts. Hence, the two papers address different institutional questions and need not deliver the same welfare conclusion.
@@ 2.1 Environment
We consider a risk-neutral planner allocating transfers across a continuum of risk-averse individuals indexed by birth cohorts. Individuals survive to the next period with probability $\alpha\in(0,1)$ and discount at rate $r\in(0,1)$; their effective discount factor is $\beta=\alpha r$. In each period, a newborn cohort enters so that the cross-section is stationary. [^3] A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game. Each individual draws an i.i.d. income shock $e_{t}\in E:=\{e^{1},e^{2},\dots,e^{M}\}\subset\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$. The planner commits to a recursive transfer mechanism $\mathcal{M}$ mapping the current promised $\lambda_{t}>0$ into a transfer $\tau_{t}$ and next period’s promise $\lambda_{t+1}$. The single state variable is the promised utility $\overline{v}_{1}(\lambda)$, which is strictly increasing in $\lambda$. We use $\lambda_{t}$ as a dual representation of promised utility $w_{t}=\overline{v}_{1}(\lambda_{t})$; all results could equivalently be written in $w_{t}$.
@@ 2.2 Symmetric-information benchmark
where $\overline{C}>0$ is an exogenous upper bound on the present value of net subsidies, $\mathbb{E}_{0}$ is the conditional expectation given the information in period 0, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period. The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above. Boundedness of $u$ is used later in the proof of Theorem 2 to show that the promised-utility martingale is bounded and hence convergent.
@@
$J(\lambda_{t},t,\tau)=(1-\beta)\mathbb{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{k}))\right]\quad\text{for }\tau\in\mathcal{T}.$ (2)
@@ 2.3 Private information and mechanism design
Definition 1 (Sequential incentive compatibility).
Given the current promise $\lambda$, truthful reporting maximizes the continuation value. Equivalently,
@@
Definition 2 (Sequentially efficient mechanism).
Fix a current state $(\lambda,e)$. Let $\mathcal{C}(\lambda,e)$ be the set of feasible sequentially incentive-compatible continuation mechanisms from that state. For any $\Gamma\in\mathcal{C}(\lambda,e)$, let
$V_{1}^{\Gamma}(\lambda,e)$
denote the individual’s expected discounted continuation utility, and let
$V_{2}^{\Gamma}(\lambda,e)$
denote the planner’s expected discounted continuation payoff (equivalently, minus the expected discounted transfer cost). We say that $\Gamma^{\prime}$ Pareto dominates $\Gamma$ at $(\lambda,e)$ if
$V_{1}^{\Gamma^{\prime}}(\lambda,e)\geq V_{1}^{\Gamma}(\lambda,e),\quad V_{2}^{\Gamma^{\prime}}(\lambda,e)\geq V_{2}^{\Gamma}(\lambda,e),$
with at least one strict inequality. A mechanism $\Gamma$ is sequentially efficient at $(\lambda,e)$ if it is sequentially incentive compatible and is not Pareto dominated at $(\lambda,e)$ by any other sequentially incentive-compatible continuation mechanism.
Remark 2.
In this paper, Pareto comparisons are made over the continuation payoff pair $(V_{1},V_{2})$, conditional on the current state $(\lambda,e)$. Accordingly, for any fixed state $(\lambda,e)$, any solution to problem (1) yields a Pareto-efficient continuation payoff pair in this sense.
@@
To construct a continuation promise satisfying both (3) and (4), define the deviation of the planner’s continuation payoff from its average by
$\Delta v_{2}(\lambda,e):=v_{2}(\lambda,e)-\overline{v}_{2}(\lambda).$
Since $v_{2}$ is the discounted value of $\tau$, a larger value of $\Delta v_{2}(\lambda,e)$ corresponds to a smaller current transfer (or a larger current contribution). We choose $\lambda^{\prime}(e\mid\lambda)$ so that this deviation is exactly offset by a change in promised continuation utility:
$\beta\lambda\bigl(v_{1}(\lambda^{\prime}(e\mid\lambda))-v_{1}(\lambda)\bigr)=\Delta v_{2}(\lambda,e).$
Equivalently,
$\lambda\Bigl[(1-\beta)u(e+\tau^{*}(e\mid\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e\mid\lambda))\Bigr]$
$=\lambda\Bigl[(1-\beta)u(e+\tau^{*}(e\mid\lambda))+\beta\overline{v}_{1}(\lambda)\Bigr]+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda).$
Hence, the continuation payoff under truthful reporting coincides with the weighted-sum problem used in (1). Moreover, since
$E[\Delta v_{2}(\lambda,e)\mid\lambda]=0,$
taking expectations yields the promise-keeping condition.
The remaining issue is feasibility. Since $\overline{v}_{1}$ is bounded above by
$S:=\sup_{x\in\mathbb{R}_{+}}u(x),$
the inverse $\overline{v}_{1}^{-1}$ is well defined only if the target continuation utility remains below $S$. If
$\overline{v}_{1}(\lambda)+\frac{1}{\beta\lambda}\bigl(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda)\bigr)<S,$
where $e^{M}$ is the maximal income realization, then $\lambda^{\prime}(e\mid\lambda)$ is well defined for every $e\in E$, and we may define
$\lambda^{\prime}(e\mid\lambda)=\overline{v}_{1}^{-1}\!\left(\overline{v}_{1}(\lambda)+\frac{1}{\beta\lambda}\bigl(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)\bigr)\right).$ (5)
@@ 3 Risk-shifting mechanism
Interpretation. Assumption 1 compares Arrow-Pratt absolute risk aversion at the contract consumption level with a scaled income spread. A convenient class of examples is given by utility functions with nonincreasing absolute risk aversion, since in that case the condition is easiest to verify at low consumption levels. In Section 4 our numerical specification
$u(c)=\frac{(c+1)^{1+\gamma}}{1+\gamma},\qquad\gamma=-3,$
has
$\frac{u^{\prime\prime}(c)}{u^{\prime}(c)}=\frac{3}{c+1},$
so absolute risk aversion is decreasing.
We define risk-shifting mechanism as follows.
Definition 3.
(Risk-shifting mechanism) A risk-shifting mechanism $\mathcal{M}$ is a sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ and transfers $\tau_{t}(e_{t})=\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ such that $\lambda_{t+1}$ is given by (6).
@@
Under Assumption 1, the feasibility correction never binds, so current consumption is constant at
$c_{t}=(u^{\prime})^{-1}(\lambda_{t}^{-1}),$
and the realized shock affects only the next-period promise. If the feasibility correction (6) binds, then agents whose unconstrained continuation promise would exceed the maximal feasible utility level receive the capped continuation promise instead. Since their contribution is capped at the threshold level, they retain part of the current income shock in current consumption. Hence, full within-period insurance fails for those realizations, but the promise-keeping condition holds weakly in discounted lifetime utility.
Remark 4.
If utility were unbounded above, the feasibility issue behind Assumption 1 would disappear, but the martingale-convergence argument would then require a different integrability condition.
@@ 3.1 Individual level: bounded promised utilities
This subsection shows that the sequence of promised utilities forms a martingale and therefore converges, so immiseration cannot occur.
@@
The promised utility process $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ is a martingale and converges to an integrable random variable a.s. (no immiseration).
@@
We have for $n\in\mathbb{N}$ $\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e. $[P]$. Hence, $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale, and promised lifetime utility is preserved in conditional expectation over time. Realized promised utility may vary across histories, but it exhibits no systematic drift. Since the process is bounded, it converges almost surely to an integrable random variable, so immiseration does not occur. In Thomas and Worrall [1990], the state variable (promised utility) $V^{t}$ converges to -$\infty$, whereas the martingale object is the lender’s marginal value process $U^{\prime}_{*}(V^{t})$.
@@ 3.2 Planner’s budget and sustainability
Proposition 1 (second-order sufficient condition for budget balance).
@@
Proof sketch. The key step is to show that the second-order approximation of the average-transfer increment $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is decreasing in $n$. For a fixed current promise $\lambda_{n-1}$ and realization $e$, define
$\Delta_{n-1}:=\overline{v}_{1}(\lambda_{n}(e\mid\lambda_{n-1}))-\overline{v}_{1}(\lambda_{n-1})=\frac{1-\beta}{\beta\lambda_{n-1}}(e-E[e_{n}]).$
We expand
$u^{-1}(\overline{v}_{1}(\lambda_{n}(e\mid\lambda_{n-1})))-u^{-1}(\overline{v}_{1}(\lambda_{n-1}))$
around $\overline{v}_{1}(\lambda_{n-1})$. As shown in Appendix C, a third-order Taylor expansion yields a second-order approximation for $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ after taking expectations. In the symmetric case, the odd-order term cancels in expectation, so the leading term is the second-order term. Assumptions (ii) and (iii), together with Jensen’s inequality, imply that this second-order term is decreasing in $n$. Hence the second-order approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\geq 1}$ is decreasing.
Therefore,
$\overline{\tau}_{n}-\overline{\tau}_{n-1}\leq\overline{\tau}_{1}-\overline{\tau}_{0}\qquad\text{(under the second-order approximation),}$
and substituting this bound into
shows that condition (iv) is sufficient for $g_{t}(\lambda_{0})\geq 0$ for all $t\in\mathbb{N}$. See Appendix C for the complete proof.
@@ 4 Numerical examples
An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Proposition 1 is given by the second-order polynomial approximation as follows.
@@
The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Proposition 1 holds with the second-order polynomial approximation if the initial promise $\lambda_{0}$ satisfies
@@
We illustrate the mechanism with a Laurent polynomial utility, a symmetric discrete income distribution, and the benchmark parameter $r=\alpha=0.93$.
@@
The values follow Corollary 1: in this calibration the lower bound $\lambda_{CE}=1/u^{\prime}\bigl(u^{-1}(E[u(e_{t})])\bigr)=49.608$ ensures higher lifetime utility than autarky across all cohorts, and the upper bound $133$ is set to satisfy (9).
@@
The series of sample means of transfer balances $\{g_{t}(\lambda_{0})\}_{t=0}^{1000}$ remains non-negative, and is consistent with Proposition 1.